Base Year Appie Bananas 7000 bunches$4 per bunch 1 000 bags place)
Answer #1

1) Nominal GDP in base year = Price of current year x Quantity of current year

= 1 x 3000 + 4 x 7000 + 4 x 8000 = 3000 + 28000 + 32000 = 63000

2) Nominal GDP in current year = 2 x 4000 + 3 x 15000 + 6 x 32000 = 8000 + 45000 + 192000 = 245000

3) % increase in Nominal GDP = (Nominal GDP in current year - Nominal GDP in base year)/Nominal GDP in base year X 100

= (245000 - 63000)/63000 x 100 = 288.9%

4) Real GDP in base year = Base year price x Base year quantity = 1 x 3000 + 4 x 7000 + 4 x 8000 = 3000 + 28000 + 32000 = 63000

5) Real GDP in current year = Base year price x Current year quantity = 1 x 4000 + 4 x 15000 + 4 x 32000

= 4000 + 60000 + 128000 = 192000

6) % increase in real GDP = (192000 - 63000)/63000 x 100 = 204.8%

7) GDP deflator for base year = Nominal GDP/Real GDP x 100 = 63000/63000 x 100 = 100

8) GDP deflator for current year = 245000/192000 x 100 = 127.6

9) Inflation rate = [(GDP deflator in current year - GDP deflator in base year) / GDP deflator in base year] X 100

= (127.6 - 100)/100 x 100 = 27.6%

10) Provide drop down options